A chemist has two solutions of HCL. One has a 40% concentration and the other has a 25% concentration.
How many liters of each solution must be mixed to obtain 124 liters of a 35% solution?
___L of the 40% solution and _____ liters of the 25% solution must be mixed to obtain a 35% solution of HCL
(Round to the nearest tenth, if necessary)

Answer by stanbon(75887)

A chemist has two solutions of HCL. One has a 40% concentration and the other has a 25% concentration.
How many liters of each solution must be mixed to obtain 124 liters of a 35% solution?
___L of the 40% solution and _____ liters of the 25% solution must be mixed to obtain a 35% solution of HCL
-------
Equation:
acid + acid = acid
0.40x + 0.25(124-x) = 0.35*124
-----
40x + 25*124 - 25x = 35*124
15x = 10*124
x = (2/3)124
x = 82 liters (amt. of 40% solution needed)
124-x = 42 liters (amt. of 25% solution needed)
==============
Cheers,
Stan H.
==============